Gas Density 3. What is the density of dinitrogen tetroxide gas at 15°C and 835 torr?
Answer #1

Pressure, P= 835 torr

We will be working in SI units.

1 torr is approximately 133.322 Pa. Thus, P= 835 torr = 835 *133.322 Pa = 111324 Pa

Temperature, T= 15 oC = 15+273 K = 298 K

Molar Mass of N2O4 = 92.011 g/mol = 92.011g/mol * (0.001 kg/g) = 0.092011 kg/mol

Value of R that is the gas constant is 8.314 J/mol-K

We will assume dinitrogen tetroxide to be an ideal gas which follows the equation.

PM = pRT

Here, \rho is the density. Putting in the values, we have,

PM = PRT (111324J/mº)(0.092011kg/mol) = ()(8.314J/mol – K)(298K) p=4.134kg/m3